OpenStudy (anonymous):

two fixed charges +q AND +4q are kept at a distance r from eachother. at what distance between the charges ,a third charge q be placed to keep them in equilibrium.

OpenStudy (anonymous):



OpenStudy (anonymous):

plz explain more

OpenStudy (anonymous):

Let a charge Q be kept at a distance of "x" from the charge 4q. All the three charges should be in eqbm. Applying eqbm condition to the Q charge k(4q)Q/x^2=kQq/(r-x)^2 Solving this we get x=2r/3 Hence charge Q will remain in eqbm if it is kept at 2r/3 from 4q no matter how large Q might be. But in order for the other charges to remain in eqbm we have to find the value of Q. q will repel 4q. So obviously Q should be negative.Now we find the magnitude of Q by applying eqbm on 4q (u may use q also). kQ(4q)/(2r/3)^2=4kqq/r^2 we will get Q=4q/9. *****USE OF q IN APPLYING EQBM WILL ALSO YIELD THE SAME VALUE OF "Q"*****

OpenStudy (anonymous):

thanx